OpenStudy (anonymous):

In a square, the sum of the lengths of three sides is 48 and in a rectangle the difference between a length and a width is 7 in. If a width of the rectangle is the same as a side of the square, how long is a length of the rectangle? Answer 17 23 32 37

jhonyy9 (jhonyy9):

square - sum of 3 sides =48 on side =16 rectangle x-y=7 - the width is the same as side of square =16 - the lenght of the recatangle will be 16+7=23